how many terms has a G.P whose second term is 1/2 and the common ratio and the last term are 1/2 and 1/120 respectively
Answer by stanbon(75887)

how many terms has a G.P whose second term is 1/2 and the common ratio and the last term are 1/2 and 1/120 respectively
-------------------
a(2) = ar = 1/2
Since r = 1/2, a = 1
-------
Last term = a(n) = ar^(n-1) = 1*(1/2)^(n-1) = 1/120
----
Solve:: (1/2)^(n-1) = 1/120
---
(n-1)log(1/2) = log(1/120)
---
n-1 = log(1/120)/log(1/2) = 6.90
n = 7.90
----
Comment:: That answer has to be a whole number.
Did you post the problem correctly?
--------------
Cheers,
Stan H.
=================